If you are finding the line perpendicular to x + 3y = 5, what is the slope of the line?


Found 2 solutions by solver91311, haileytucki:
Answer by solver91311(24713)



Solve your given equation for y in terms of everything else. The resulting coefficient on x will be the slope of the given line. The slope of the perpendicular will be the negative reciprocal of the slope of the given line.



John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism



Answer by haileytucki(390)

x+3y=5
Since x does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting x from both sides.
3y=-x+5
Divide each term in the equation by 3.
(3y)/(3)=-(x)/(3)+(5)/(3)
Simplify the left-hand side of the equation by canceling the common factors.
y=-(x)/(3)+(5)/(3)
To find the slope and y intercept, use the y=mx+b formula where m=slope and b is the y intercept.
y=mx+b
Using the y=mx+b formula, m=-(1)/(3) and b=(5)/(3).
m=-(1)/(3), b=(5)/(3)